### v1

# Primordial black holes with anisotropic hair

Abstract
A mechanism for generating anisotropic enhancements of the curvature perturbation through a vector field is proposed. We find that when the mixing between the inflaton perturbation and the vector-field perturbation is sufficiently strong in the anisotropic inflation, the power spectrum becomes dominated by anisotropic constant modes. This suggests that statistical anisotropy in primordial black hole (PBH) formation may be inevitable if inflation undergoes an anisotropic inflationary phase. Our findings offer a novel approach to probe vector fields during inflation and to test the cosmic no-hair conjecture.
preprint: APS/123-QED
I Introduction
The small-scale physics of inflation remains poorly understood. The cosmic microwave background(CMB) observation only constrains the first few e-folds of inflation—yielding an almost scale-invariant, Gaussian and adiabatic curvature perturbation on large scales($k\lesssim 1\ \text{Mpc}^{-1}$) [1, 2]. On small scales, however, the curvature perturbation may be responsible for the creation of PBH with an abundance compatible with dark matter [3, 4, 5, 6, 7]. Such PBH production requires the power spectrum to exhibit a peak(which strongly depends on $k$) on small scales with an amplitude at least $10^{7}$ times larger than that of the CMB. The enhancement of the amplitude of the curvature perturbation can be generated in several mechanisms [8, 9, 10, 11, 12, 13, 14].
In this Letter, we explore the possibility of large anisotropic curvature perturbation sourced by a vector field, and examine their imprint on PBH production when the vector field is amplified through the kinetic coupling $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$, which is originally motivated by supergravity [15]. A broad class of coupling functions permits the so-called anisotropic inflation, characterized by a constant electric field during inflation [16, 17, 18, 19, 20, 21]. The key parameter of the phenomenology in anisotropic inflation is $\mathcal{I}$(defined explicitly in the next section), which characterizes the mixing between the inflaton perturbation and the vector-field perturbation. This mixing term induces anisotropy in the power spectrum of the curvature perturbation $\mathcal{R}$. Previous work derived this anisotropic contribution as $\delta\mathcal{P}_{\mathcal{R}}/\mathcal{P}_{\mathcal{R},0}=16\mathcal{I}^{2}N_{k}^{2}\sin^{2}\theta$ [22, 23, 24, 25, 26, 27, 28], where $N_{k}\sim 60$ represents the e-folding number of the mode elapsed since the horizon-crossing, and $\theta$ is the angle between the vector field and the wavenumber. The factor $N_{k}^{2}$ arises from the cumulative effect of vector-field perturbations acting as a source on superhorizon scales, driving the evolution of $\mathcal{R}$. It seems that we can take $\mathcal{I}\gg 1$ of this result to obtain a large enhancement of the power spectrum. However, it has been shown that the in-in perturbation theory breaks down for $\mathcal{I}\gtrsim\mathcal{O}(0.1)$  [^1] Although [55, 46] studied the isotropic configuration of a triad of vector fields, the conclusion should also hold for the case of a single vector field. The strong mixing of perturbations has been discussed recently in the context of multi-field inflation [30, 31, 32, 33, 34, 35, 36, 37]. One may concern about applying similar treatments to vector fields.
The purpose of this Letter is to investigate the regime $\mathcal{I}\gg 1$ of anisotropic inflation  [^2] The isotropic case of a triad of vectors has been studied in [54, 21, 46]. Here we focus on the anisotropic one. We find that in this regime, the dominant contribution to $\mathcal{R}$ arises not from cumulative-effect modes but from constant modes when $\mathcal{I}\sin\theta\gg 1$. To elucidate this mechanism, we will be in the perspective of the effective field theory(EFT) by integrating out heavy modes. Our analysis reveals that the power spectrum exhibits exponential enhancement:
$\Delta_{\mathcal{R}}\equiv\frac{\mathcal{P}_{\mathcal{R}}}{\mathcal{P}_{\mathcal{R},0}}\sim e^{4\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}$ (1)
for $\sin^{2}\theta>1/2$, where $\mathcal{P}_{\mathcal{R},0}\equiv\mathcal{P}_{\mathcal{R}}(\mathcal{I}=0)$. Only modes with $\theta\sim\pi/2$ are sufficiently enhanced, leading to a statistically anisotropic power spectrum. If inflation experiences a transient anisotropic phase, such anisotropic peaks will inevitably imprint on the power spectrum, with potential cosmological implications.
II Anisotropic inflation
We begin by reviewing anisotropic inflation. Consider a model coupling an inflaton field to a vector field through the action
$S=\int d^{4}x\sqrt{-g}\bigg{[}$ $\frac{M_{\text{pl}}^{2}}{2}R-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)$
$-\frac{f^{2}(\phi)}{4}F_{\mu\nu}F^{\mu\nu}\bigg{]},$ (2)
where the kinetic function $f(\phi)$ governs the coupling between the inflaton and the vector field. The equation of motion of $A_{i}(t)$ can be solved as $\dot{A}_{i}(t)\propto f^{-2}a^{-1}$, with the overdot denoting the derivative with respect to the cosmological time $t$. The energy densities of the electric and magnetic fields are given by [39]
$\rho_{E}\sim\frac{f^{2}}{a^{2}}\dot{A}_{i}^{2},\ \ \ \ \ \ \ \rho_{B}\sim\frac{f^{2}}{2a^{4}}\left(\partial_{i}A_{j}-\partial_{j}A_{i}\right)^{2}.$ (3)
We are interested in a constant electric field that requires $f\sim a^{-2}$. For a homogeneous vector field($\rho_{B}=0$), the universe remains spatially homogeneous. However, the non-zero $\rho_{E}$ maintains an anisotropic spacetime during the evolution of the universe. Without loss of generality, adopting the ansatz $A_{\mu}=(0,A(t),0,0)$, the metric of spacetime can be represented as a Bianchi type I form $ds^{2}=-a^{2}(\eta)\left(-d\eta^{2}+dx^{2}\right)+b^{2}(\eta)\left(dy^{2}+dz^{2}\right)$, where $\eta$ is the conformal time $d\eta=dt/a$. If the energy density of the vector field remains sub-dominated compared with inflaton $\rho_{A}\ll V$, the anisotropy of spacetime is always of the order of slow-roll parameters [22], i.e. $aH\simeq(-\eta)^{-1-\epsilon_{H}}$ and $bH\simeq(-\eta)^{-1-\epsilon_{H}-\epsilon_{A}}$ so that $b^{\prime}/b-a^{\prime}/a\simeq aH\epsilon_{A}$, where the prime denotes the derivative with respect to conformal time $\eta$.
During the anisotropic phase, the potential of inflaton still dominates the inflation while a portion of its energy transforms to the vector field. In this scenario, both the rolling of the inflaton and the vector field are made responsible for the deviation of the de Sitter space, i.e., $\epsilon_{H}=\epsilon_{\phi}+\epsilon_{A}$, where $\epsilon_{H}\equiv-\dot{H}/H^{2},\epsilon_{\phi}\equiv\dot{\phi}^{2}/2M_{\text{pl}}^{2}H^{2},\epsilon_{A}\equiv f^{2}\dot{A}^{2}/3a^{2}M_{\text{pl}}^{2}H^{2}$. It is convenient to define a parameter:
$\mathcal{I}\equiv\sqrt{\frac{3\epsilon_{A}}{2\epsilon_{\phi}}},$ (4)
which characterizes the ratio between two rolling velocity of the fields.
[…]
III Strong mixing of perturbations
The vector field breaks the 3-d rotational symmetry of the universe to a 2-d rotational one in the $yz$ plane. This symmetry breaking complicates the metric perturbations, with the complete quadratic action having been thoroughly analyzed in previous works [22, 23, 24, 25, 26, 27, 28].
However, it is known that the dominated contributions to the curvature perturbation arise only from matter perturbations in the spatially-flat gauge, while the interactions between matter and metric are slow-roll suppressed. Consequently, we may safely neglect gravitational back-reaction in the following. The matter perturbations are given by
$\phi=\phi_{0}+\varphi,\ \ \ \ \ A_{\mu}^{(S)}=(\delta A_{0},A+\delta A_{1},0,0).$ (5)
Here we consider only the scalar parts of the perturbations of $A_{\mu}$ and have removed d.o.f by fixing the U(1) symmetry. Without loss of generality, we align the wavenumber as $\bm{k}=(k_{1},k_{2},0)$, where $k_{1}=k\cos\theta$ and $k_{2}=(b/a)k\sin\theta$, and $\theta$ being the angle between the vector field and $k$. After eliminating the non-dynamical d.o.f. $\delta A_{0}$ and introducing a new variable $\mathcal{A}\equiv(f/a)\sin\theta\delta A_{1}$, the dominant quadratic action is given by
$\mathcal{L}_{2}\simeq\frac{a^{3}}{2}\Bigg{[}$ $|\dot{\varphi}|^{2}-\left(\frac{k^{2}}{a^{2}}+8\mathcal{I}^{2}H^{2}\cos{2\theta}\right)|\varphi|^{2}+|\dot{\mathcal{A}}|^{2}$
$-\frac{k^{2}}{a^{2}}|\mathcal{A}|^{2}-4H\mathcal{I}\sin\theta\Big{(}\varphi^{*}\dot{\mathcal{A}}+c.c\Big{)}$
$-12H^{2}\mathcal{I}\sin\theta\Big{(}\varphi^{*}\mathcal{A}+c.c.\Big{)}\Bigg{]}.$ (6)
Here, we have assumed $\dot{\phi}<0$, $\dot{A}>0$, and treat $\mathcal{I}$ as a constant to leading-order of the slow-roll approximation. Spacetime anisotropy, being slow-roll suppressed, is also neglected.
The last two terms in (III) represent the mixing between $\varphi$ and $\mathcal{A}$, which is controlled by $\mathcal{I}\sin\theta$. We are interested in the strong mixing $\mathcal{I}\sin\theta\gg 1$, which is the subject of the rest of this paper. One can immediately note that in this regime the mass $m^{2}_{\varphi}\equiv 8\mathcal{I}^{2}H^{2}\cos{2\theta}$ can become heavy for some angle values $\theta$. If this is the case, there is a hierarchy between two frequencies $\omega_{\pm}^{2}=p^{2}+4\mathcal{I}^{2}H^{2}\pm 4\mathcal{I}H(\mathcal{I}^{2}H^{2}+p^{2}\sin^{2}\theta)^{1/2}$ of modes on sub-horizon scales [40]. When the momentum of the modes redshifts to $p\ll|m_{\varphi}|$, the high-frequency mode $\omega_{+}$ can be integrated out, yielding a low-energy effective theory for the light modes [41, 42, 43].
Specifically, To obtain the leading-order solution for $\varphi$, we expand the propagator $1/(m^{2}_{\varphi}-\square)=1/m^{2}_{\varphi}+\square/m^{4}_{\varphi}+\cdots$ in the equation of motion. Then we can solve this leading solution as
$\varphi_{\text{LO}}=-\frac{\sin\theta\big{(}\dot{\mathcal{A}}+3H\mathcal{A}\big{)}}{2\mathcal{I}H\cos{2\theta}}$ (7)
for $|m_{\varphi}^{2}|\gg H^{2}$. The heavy d.o.f can be integrated out by substituting this solution into the quadratic action. Then the leading-order action of the effective theory of $\mathcal{A}$ is given by
$\mathcal{L}_{2,\text{LO}}\simeq\frac{a^{3}}{2c_{s}^{2}(\theta)}\left[|\dot{\mathcal{A}}|^{2}-\frac{c_{s}^{2}(\theta)k^{2}}{a^{2}}|\mathcal{A}|^{2}\right],$ (8)
where the sound speed is given by
$c_{s}^{2}(\theta)=\cos{2\theta}.$ (9)
This is an effective action of a massless perturbation with sound speed $c_{s}$. Remarkably, $c_{s}^{2}$ becomes imaginary for $\theta\in(\pi/4,3\pi/4)\cup(5\pi/4,7\pi/4)$, leading to exponentially growing solutions before the horizon crossing  [^3] Recently, the multi-field inflation with strongly non-geodesic attractors in the field space has been extensively explored [57, 58]. The effective theory of inflationary dynamics is a massless mode with imaginary sound speed [30, 31, 32, 33, 34, 35]. It can also provide the mechanism for generating PBH in the multi-filed scenario [36, 37]. Consequently, the power spectrum of $\mathcal{A}$ is scale-invariant(of leading order) and enhanced on large scales. This enhancement is anisotropic due to the angle-dependent sound speed $c_{s}(\theta)$. On the other hand, if $\sin\theta\simeq 0$, $\varphi$ and $\mathcal{A}$ are decoupled and $\varphi$ is also heavy with a real mass and decay away on the superhorizon for any value of large $\mathcal{I}$.
The enhanced solution for $\mathcal{A}$ on the superhorizon can be found by solving the coupled system inside the horizon [32, 45, 46]. Here we are only interested in the enhanced solution, where $\sin^{2}{\theta}>1/2$. The conjugate momentum of the perturbations is given by
$\pi_{\varphi}=\frac{a^{3}}{2}\dot{\varphi},\ \ \ \ \ \pi_{\mathcal{A}}=\frac{a^{3}}{2}\left(\dot{\mathcal{A}}-4\mathcal{I}\sin{\theta}H\varphi\right).$ (10)
After defining $p_{\varphi}\equiv 2\pi_{\varphi}/a^{3}H$, $p_{\mathcal{A}}\equiv 2\pi_{\mathcal{A}}/a^{3}H$, $x\equiv-k\eta$  and using the e-folding number as the time variable, the coupled system (III) gives $X_{,N}=J_{0}X+J_{1}X$, where $X=(\mathcal{A},p_{\mathcal{A}},\varphi,p_{\varphi})^{T}$,
$J_{0}=$ $\left(\ \begin{matrix}0&1&4\mathcal{I}\sin{\theta}&0&\\ -x^{2}&0&0&0&\\ 0&0&0&0&\\ 0&-4\mathcal{I}\sin{\theta}&-(x^{2}+8\mathcal{I}^{2})&0&\end{matrix}\right),$
$J_{1}=$ $\left(\ \begin{matrix}0&0&0&0&\\ 0&-3&-12\mathcal{I}\sin{\theta}&0&\\ 0&0&0&0&\\ -12\mathcal{I}\sin{\theta}&0&0&-3&\end{matrix}\right).$ (11)
Considering the Hubble term as a small correction to the system on subhorizon, the matrix $J_{0}$ is of 00-$th$ order while $J_{1}$ is $1$-$st$ order. Near the horizon crossing, the exponential growth mode of $\mathcal{A}$ emerges from the imaginary frequency $\omega_{-}$ of the coupled system when $p<|m_{\varphi}|$, which yields the enhanced solution at the limit $-k\eta\to 0$ [47],
$\mathcal{A}\simeq\left|2-4\sin^{2}\theta\right|^{1/4}e^{2\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}.$ (12)
The enhanced factor is, as expected from the sound speed of the EFT, dependent on the angle $\theta$.
One thing to mention is that even on superhorizon scales, $\varphi$ and $\mathcal{A}$ remain coupled for $\sin\theta\neq 0$. This can be realized from solution (7) that when $\dot{\mathcal{A}}=0$ on the superhorizon, the two perturbations are still totally relevant, i.e.,
$\varphi_{\text{LO}}=-\frac{3\sin\theta}{2\mathcal{I}\cos{2\theta}}\mathcal{A},\ \ \ \ \left(-k\eta\ll 1\right)$ (13)
We found that in this limit, the perturbation $\mathcal{A}$ is dominates over the perturbation $\varphi$. This persistent coupling implies that if anisotropic inflation with $\mathcal{I}\gg 1$ occurs after the time that the CMB modes exit the horizon, these large-scale modes can still be kicked by the vector field.
IV Seeding PBH from a vector field
IV.1 Features of the power spectrum
The PBH can be formed if the curvature perturbation generated on small scales during inflation is enhanced. For an anisotropic inflationary universe, the uniform-energy curvature perturbation we can use in the spatially-flat gauge is
$\zeta\simeq-H\frac{\delta\rho}{\dot{\rho}}.$ (14)
One can directly compute the perturbation of the total energy density of matter to get the curvature perturbation. However, alternatively, here we compute a variable:
$\mathcal{R}=H\frac{ik^{a}}{k^{2}}V_{a}$ (15)
in the spatially-flat gauge, where $a=1,2,3$ and the total velocity $V_{a}\equiv(\rho+P)^{-1}T^{t}_{\ a}$. This reduces to the standard comoving curvature perturbation $\mathcal{R}_{\text{iso}}=-H(\rho+P)u$ in isotropic backgrounds, where $u$ is the scalar velocity of fluids $v_{a}=\partial_{a}u+u_{a}$. One can find that in our configuration of anisotropic inflation, $\mathcal{R}=-\zeta$ holds on superhorizon scales up to slow-roll corrections from spacetime anisotropy [47]. We therefore use $\mathcal{R}$ to compute the power spectrum, consistent with our treatment of neglecting slow-roll suppressed terms.
The curvature perturbation of the anisotropic inflation is then given by
$\mathcal{R}=\frac{1}{\sqrt{2\epsilon_{H}}M_{\text{pl}}H}\frac{-\varphi+\mathcal{I}\sin\theta\mathcal{A}}{\sqrt{1+2\mathcal{I}^{2}/3}}.$ (16)
In the EFT theory of $\mathcal{A}$ (13), the dominant contribution to the curvature perturbation is from the perturbation of the vector field $\mathcal{A}$ due to $\mathcal{R}_{\mathcal{A}}/\mathcal{R}_{\varphi}\sim\mathcal{I}^{2}\gg 1$. If we assume that $\epsilon_{H}$ is a constant during inflation, we have
$\mathcal{R}\propto\sqrt{\frac{3}{2}}\sin{\theta}\mathcal{A}.$ (17)
The Fourier modes of $\mathcal{R}$ can be decomposed in terms of two sets of operators since we have two perturbations: ~$\tilde{\mathcal{R}}(\bm{k},\eta)=\sum_{\alpha=\pm}\left[\mathcal{R_{\alpha}}(k,\eta)a_{\alpha}(\bm{k})+\mathcal{R^{*}_{\alpha}}(k,\eta)a_{\alpha}^{\dagger}(-\bm{k})\right]$, where $a_{\alpha}^{\dagger}(\bm{k})$ and $a_{\alpha}(\bm{k})$ are creation and annihilation operators that satisfy the commutation relation $[a_{\alpha}(\bm{k}),a_{\beta}^{\dagger}(\bm{k^{\prime}})]=(2\pi)^{3}\delta_{\alpha\beta}\delta^{(3)}(\bm{k}-\bm{k^{\prime}})$. The dimensionless power spectrum of the curvature perturbations is then computed by
$\mathcal{P}_{\mathcal{R}}(k)=\frac{k^{3}}{2\pi^{2}}\sum_{\alpha=\pm}|\mathcal{R}_{\alpha}(k)|^{2}.$ (18)
Here we consider a transient anisotropic phase, i.e., $\mathcal{I}=\mathcal{I}_{m}\gg 1$ in a period $\Delta N\gtrsim\log\mathcal{I}$ near $N_{f}>N_{\text{CMB}}$, to see how a time-dependent $\mathcal{I}(\eta)$ can result in an anisotropic peak of the curvature power spectrum on small scales  [^4] We assume that at time $N\ll N_{f}$ and $N\gg N_{f}$, inflation is in the anisotropic phase with $\mathcal{I}\ll 1$ so that the gauge field $\mathcal{A}$ is approximately a massless field.. In this case, only modes with wavenumber $k\sim k_{f}\equiv a(N_{f})H$ can experience exponential growth when crossing the horizon. The power spectrum of the enhanced curvature perturbation near $k\sim k_{f}$ can be estimated as
$\Delta_{\mathcal{R}}\simeq\left(\frac{3}{2}\sin^{2}{\theta}\right)\mathcal{P}_{\mathcal{A}}(k),$ (19)
where $\mathcal{P}_{\mathcal{A}}=\langle\mathcal{A}\mathcal{A}\rangle$. As is known from the solution of $\mathcal{A}$ (12), the power spectrum depends exponentially on $\sin\theta$. The maximum enhancement is at $\theta=\pi/2$. Even a slight deviation $\Delta\theta\sim\mathcal{I}^{-1}$ from $\pi/2$ leads to a drastic suppression of the peak of the power spectrum.
On the other hand, we define the mass-crossing time $k/a(\tilde{N})=|m_{\varphi}(\tilde{N})|$ for any $k$ mode. Because most of the growth occurs at the early time [32, 45, 37]. The location of the peak $k_{p}$ is determined by $\tilde{N}\sim N_{f}$, i.e. the mode that crosses the mass-horizon at time $N_{f}$,
$k_{p}(\theta)\sim\frac{|m_{\varphi}(\theta)|}{H}\cdot k_{f},$ (20)
which means that the location of the peaks is also anisotropic.
Modes with $k\gg k_{f}$ experience nothing because the vector field has been turned off well before horizon-crossing. On the other hand, large-scale modes $k\gg k_{f}$ are still almost scale-invariant but are kicked by the vector, which results in a shift deformation. The amplitudes of the shift obviously depend on the angle $\theta$, and the large-scale power spectrum is significantly anisotropy. This is inconsistent with the CMB observation. However, this work presents one minimal mechanism for generating anisotropic peaks of the power spectrum at small scales. The large-scale power spectrum may originate from alternative mechanisms, such as the curvaton scenario, depending upon the specific model framework. we provide a toy example in the following.
IV.2 A toy example
We consider a model containing two scalar fields and a vector field $\mathcal{L}=\mathcal{L}_{\phi\text{-}A}+\mathcal{L}_{\psi}$, where
$\mathcal{L}_{\phi\text{-}A}$ $=\sqrt{-g}\bigg{[}-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)-\frac{f^{2}(\phi)}{4}F_{\mu\nu}F^{\mu\nu}\bigg{]},$
$\mathcal{L}_{\psi}$ $=\sqrt{-g}\bigg{[}-\frac{1}{2}\partial^{\mu}\psi\partial_{\mu}\psi-U(\psi)\bigg{]}.$ (21)
$\mathcal{L}_{\psi}$ is only coupled to $\mathcal{L}_{\phi\text{-}A}$ through gravitational interaction. Hence, the power spectrum of the curvature perturbation of this system is given by
$\mathcal{P}_{\mathcal{R}}(k)$ $\simeq\mathcal{P}_{\psi}(k)+\mathcal{P}_{\phi\text{-}A}(k)$
$\propto\frac{r^{2}|\delta\psi|^{2}+|-\varphi+\mathcal{I}\sin{\theta\mathcal{A}}|^{2}}{r^{2}+1+2\mathcal{I}^{2}/3},$ (22)
where $r\equiv|\dot{\psi}/\dot{\phi}|$.
In the early stage of inflation, the energy density of $A$ is negligible, reducing the system to a two-field system. Assuming slow-roll conditions for both $\psi$ and $\phi$, adiabatic and entropy perturbations remain decoupled at early times [49]. Consequently, the curvature perturbation remains conserved on large scales. The ratio $r$ evolves as $r(t)=r_{\text{ini}}[1+2\mathcal{I}(t)^{2}/3]$ due to the anisotropic phase of the late time. We assume that $r_{\text{ini}}\gg 1$. In this regime, $\mathcal{P}_{\mathcal{R}}\simeq\mathcal{P}_{\psi}=(H/\dot{\psi})^{2}|\delta\psi|^{2}=2.1\times 10^{-9}$ dominates the power spectrum on large scales($k\lesssim 1\ \text{Mpc}^{-1}$). At time $N_{f}$ after the time in which the CMB modes exit the horizon, $\mathcal{I}$ becomes significant. Then the power spectrum is dominated by the $\mathcal{P}_{\phi\text{-}A}$ on scale $k_{f}$. One of the examples of this model is shown in FIG. 1. We can obviously observe the anisotropic enhancements and the anisotropic peaks(the black curve).
Figure 1: The power spectrum of $\mathcal{R}$ for our toy example (IV.2). We use the Gaussian profile $\mathcal{I}(N)=\mathcal{I}_{\text{max}}e^{-(N-N_{f})^{2}/(2\Delta N^{2})}$ with $\mathcal{I}_{\text{max}}=9$, $\Delta N=2$, $N_{f}\simeq N_{\text{CMB}}+5$ and $r_{\text{ini}}=10$. The black curve represent peaks $k_{p}(\theta)$ of $\mathcal{P}_{\mathcal{R}}$.
V Conclusion
We have shown that anisotropic inflation allows for an exponential enhancement in the power spectrum for $\mathcal{I}\gg 1$. This work presents the first fully analytical treatment of the constant, large $\mathcal{I}$ regime, revealing fundamentally different behavior from previous studies based on a small $\mathcal{I}$. Crucially, we find that superhorizon power spectrum contributions are dominated by enhanced constant modes.
The most interesting feature of the enhanced power spectrum sourced from the vector itself is the statistical anisotropy. Within the EFT framework, the light field $\mathcal{A}$ develops a $\theta$-dependent sound speed $c_{s}^{2}=\cos 2\theta$ during horizon crossing. Notably, when $c_{s}^{2}<0$, the imaginary sound speed triggers exponential growth of the light field.
This mechanism is not restricted to specific models — it can be triggered whenever the choices of $f(\phi)$ and $V(\phi)$ drive inflation into an anisotropic phase. Therefore, once inflation enters an anisotropic phase at a certain time $N_{f}$, it produces a corresponding anisotropic peak in the power spectrum at scale $k_{p}$. However, several aspects require careful consideration. We need to future study their non-Gaussianity(NG) and perturbative control. One may be concerned that the exponential enhancement of perturbations leads to an incredible enhancement on NG. However, it has been shown that the low-energy EFT with imaginary sound speed reduced from strongly non-geodesic attractors of multifield inflation generates only mild NG [32, 31, 33, 34]. Moreover, the perturbative validity is also well controlled in this attractor [31, 32, 35]. We believe that this analysis can be used in our case in this Letter.
The anisotropic features of the peak can be caught by some future observations, such as PBH production. As we know, the PBH production of a statistically anisotropic peak is still absent. It is necessary to figure out the physical details of this case. In particular, anisotropic hair may be imprinted on the initial spin of PBHs, leaving behind intriguing observational signatures [50]. Moreover, enhancements of the curvature power spectrum on small scales are generally accompanied by the production of scalar-induced gravitational waves(SIGWs). This implies that anisotropic features will also manifest as sources in the SIGW spectrum. The frequencies of SIGWs is related to $N_{f}$ as $N_{f}+\log\left(\mathcal{I}\right)\sim\log\left(f_{\text{peak}}/\text{Hz}\right)+37$ [51], which spans a broad spectrum covered by gravitational-wave experiments, such as PTA and LISA. A model-independent discussion of SIGWs sourced from a statistically anisotropic scalar spectrum has been studied in [52, 53]. We provide a specific realization of a statistically anisotropic peak in this Letter. It is interesting to explore signatures of this anisotropic hair in the SIGWs and test it in future observations.
Acknowledgements.
We would like to thank Jiro Soda for helpful discussion. This work was supported by the National Natural Science Foundation of China with the Grant No. 12375049, and Key Program of the Natural Science Foundation of Jiangxi Province under Grant No. 20232ACB201008.

### v2

@@
 The small-scale physics of inflation remains poorly understood.
-The cosmic microwave background(CMB) observation only constrains the first few e-folds of inflation—yielding an almost scale-invariant, Gaussian and adiabatic curvature perturbation on large scales($k\lesssim 1\ \text{Mpc}^{-1}$) [1, 2].
-On small scales, however, the curvature perturbation may be responsible for the creation of PBH with an abundance compatible with dark matter [3, 4, 5, 6, 7].
+The cosmic microwave background(CMB) observation only constrains the first few $e$-folds of inflation—yielding an almost scale-invariant, Gaussian and adiabatic curvature perturbation on large scales($k\lesssim 1\ \text{Mpc}^{-1}$) [Lyth:1996im, Planck:2018jri].
+On small scales, however, the curvature perturbation may be responsible for the creation of PBH with an abundance compatible with dark matter [Hawking:1971ei, Carr:1974nx, Chapline:1975ojl, Carr:2016drx, Inomata:2017okj].
 Such PBH production requires the power spectrum to exhibit a peak(which strongly depends on $k$) on small scales with an amplitude at least $10^{7}$ times larger than that of the CMB.
-The enhancement of the amplitude of the curvature perturbation can be generated in several mechanisms [8, 9, 10, 11, 12, 13, 14].
-In this Letter, we explore the possibility of large anisotropic curvature perturbation sourced by a vector field, and examine their imprint on PBH production when the vector field is amplified through the kinetic coupling $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$, which is originally motivated by supergravity [15].
-A broad class of coupling functions permits the so-called anisotropic inflation, characterized by a constant electric field during inflation [16, 17, 18, 19, 20, 21].
-The key parameter of the phenomenology in anisotropic inflation is $\mathcal{I}$(defined explicitly in the next section), which characterizes the mixing between the inflaton perturbation and the vector-field perturbation.
-This mixing term induces anisotropy in the power spectrum of the curvature perturbation $\mathcal{R}$.
-Previous work derived this anisotropic contribution as $\delta\mathcal{P}_{\mathcal{R}}/\mathcal{P}_{\mathcal{R},0}=16\mathcal{I}^{2}N_{k}^{2}\sin^{2}\theta$ [22, 23, 24, 25, 26, 27, 28], where $N_{k}\sim 60$ represents the e-folding number of the mode elapsed since the horizon-crossing, and $\theta$ is the angle between the vector field and the wavenumber.
+The enhancement of the amplitude of the curvature perturbation can be generated in several mechanisms [Ozsoy:2023ryl, Jedamzik:1999am, Bugaev:2013fya, Linde:2012bt, Kawasaki:2019hvt, Cai:2018tuh, Green:2000he].
+In this paper, we explore the possibility of large anisotropic curvature perturbation sourced by a vector field and examine their imprint on PBH production when the vector field is amplified through the kinetic coupling $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$, which is originally motivated by supergravity [Ratra:1991bn].
+A broad class of coupling functions permits the so-called anisotropic inflation, characterized by a constant electric field during inflation [Watanabe:2009ct, Maleknejad:2012fw, Kanno:2010nr, Soda:2012zm, Watanabe:2010bu, Chen:2022ccf].
+The key parameter of the phenomenology in anisotropic inflation is $\mathcal{I}$:
+$\mathcal{I}\equiv\sqrt{\frac{3\epsilon_{A}}{2\epsilon_{\phi}}},$ (1)
+which characterizes the mixing between the inflaton perturbation and the vector-field perturbation.
+Here $\epsilon_{\phi}$ and $\epsilon_{A}$ are the velocity of the inflaton and a vector field respectively.
+The mixing term induces anisotropy in the power spectrum $\mathcal{P}_{\mathcal{R}}$ of the curvature perturbation $\mathcal{R}$.
+Previous work derived this anisotropic contribution as $\delta\mathcal{P}_{\mathcal{R}}/\mathcal{P}_{\mathcal{R},0}=16\mathcal{I}^{2}N_{k}^{2}\sin^{2}\theta$ [Watanabe:2010fh, Emami:2013bk, Bartolo:2012sd, Gumrukcuoglu:2007bx, Himmetoglu:2009mk, Gumrukcuoglu:2010yc, Durrer:2024ibi], where $\mathcal{P}_{\mathcal{R},0}\equiv\mathcal{P}_{\mathcal{R}}(\mathcal{I}=0)$, $N_{k}\sim 60$ represents the $e$-folding number of the mode elapsed since the horizon-crossing, and $\theta$ is the angle between the vector field and the wave number.
 The factor $N_{k}^{2}$ arises from the cumulative effect of vector-field perturbations acting as a source on superhorizon scales, driving the evolution of $\mathcal{R}$.
 It seems that we can take $\mathcal{I}\gg 1$ of this result to obtain a large enhancement of the power spectrum.
-However, it has been shown that the in-in perturbation theory breaks down for $\mathcal{I}\gtrsim\mathcal{O}(0.1)$  [^1] Although [55, 46] studied the isotropic configuration of a triad of vector fields, the conclusion should also hold for the case of a single vector field.
-The strong mixing of perturbations has been discussed recently in the context of multi-field inflation [30, 31, 32, 33, 34, 35, 36, 37].
+However, it has been shown that the in-in perturbation theory breaks down for $\mathcal{I}\gtrsim\mathcal{O}(0.1)$  [^1] Although [Funakoshi:2012ym, Chen:2023bcz] studied the isotropic configuration of a triad of vector fields, the conclusion should also hold for the case of a single vector field.
+The strong mixing of perturbations has been discussed recently in the context of multifield inflation [Garcia-Saenz:2018ifx, Fumagalli:2019noh, Bjorkmo:2019qno, Garcia-Saenz:2019njm, Garcia-Saenz:2018vqf, Garcia-Saenz:2025jis] and is used as a new mechanics of PBH formation [Palma:2020ejf, Fumagalli:2020adf, Braglia:2020eai].
 One may concern about applying similar treatments to vector fields.
-The purpose of this Letter is to investigate the regime $\mathcal{I}\gg 1$ of anisotropic inflation  [^2] The isotropic case of a triad of vectors has been studied in [54, 21, 46].
+The purpose of this paper is to investigate the regime $\mathcal{I}\gg 1$ of anisotropic inflation  [^2] The isotropic case of a triad of vectors has been studied in [Yamamoto:2012sq, Chen:2022ccf, Chen:2023bcz].
 Here we focus on the anisotropic one.
@@
 $\Delta_{\mathcal{R}}\equiv\frac{\mathcal{P}_{\mathcal{R}}}{\mathcal{P}_{\mathcal{R},0}}\sim e^{4\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}$ (2)
-for $\sin^{2}\theta>1/2$, where $\mathcal{P}_{\mathcal{R},0}\equiv\mathcal{P}_{\mathcal{R}}(\mathcal{I}=0)$.
+for $\sin^{2}\theta>1/2$.
 Only modes with $\theta\sim\pi/2$ are sufficiently enhanced, leading to a statistically anisotropic power spectrum.
@@
 $-\frac{f^{2}(\phi)}{4}F_{\mu\nu}F^{\mu\nu}\bigg],$ (3)
-where the kinetic function $f(\phi)$ governs the coupling between the inflaton and the vector field.
-The equation of motion of $A_{i}(t)$ can be solved as $\dot{A}_{i}(t)\propto f^{-2}a^{-1}$, with the overdot denoting the derivative with respect to the cosmological time $t$.
-The energy densities of the electric and magnetic fields are given by [39]
+where $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, and the kinetic function $f(\phi)$ governs the coupling between the inflaton and the vector field.
+After gauge fixing $A_{0}=0$, the equation of motion of $A_{i}(t)$ can be solved as $\dot{A}_{i}(t)\propto f^{-2}a^{-1}$, where the overdot denoting the derivative with respect to the cosmological time $t$.
+The energy densities of the electric and magnetic fields are given by [Subramanian:2009fu]
 $\rho_{E}\sim\frac{f^{2}}{a^{2}}\dot{A}_{i}^{2},\ \ \ \ \ \ \ \rho_{B}\sim\frac{f^{2}}{2a^{4}}\left(\partial_{i}A_{j}-\partial_{j}A_{i}\right)^{2}.$ (4)
@@
 For a homogeneous vector field($\rho_{B}=0$), the Universe remains spatially homogeneous.
-However, the non-zero $\rho_{E}$ maintains an anisotropic spacetime during the evolution of the universe.
+However, nonzero $\rho_{E}$ maintains an anisotropic spacetime during the evolution of the universe.
 Without loss of generality, adopting the ansatz $A_{\mu}=(0,A(t),0,0)$, the metric of spacetime can be represented as a Bianchi type I form $ds^{2}=-a^{2}(\eta)\left(-d\eta^{2}+dx^{2}\right)+b^{2}(\eta)\left(dy^{2}+dz^{2}\right)$, where $\eta$ is the conformal time $d\eta=dt/a$.
-If the energy density of the vector field remains sub-dominated compared with inflaton $\rho_{A}\ll V$, the anisotropy of spacetime is always of the order of slow-roll parameters [22], i.e.
-$aH\simeq(-\eta)^{-1-\epsilon_{H}}$ and $bH\simeq(-\eta)^{-1-\epsilon_{H}-\epsilon_{A}}$ so that $b^{\prime}/b-a^{\prime}/a\simeq aH\epsilon_{A}$, where the prime denotes the derivative with respect to conformal time $\eta$.
-During the anisotropic phase, the potential of inflaton still dominates the inflation while a portion of its energy transforms to the vector field.
-In this scenario, both the rolling of the inflaton and the vector field are made responsible for the deviation of the de Sitter space, i.e., $\epsilon_{H}=\epsilon_{\phi}+\epsilon_{A}$, where $\epsilon_{H}\equiv-\dot{H}/H^{2},\epsilon_{\phi}\equiv\dot{\phi}^{2}/2M_{\text{pl}}^{2}H^{2},\epsilon_{A}\equiv f^{2}\dot{A}^{2}/3a^{2}M_{\text{pl}}^{2}H^{2}$.
-It is convenient to define a parameter:
-$\mathcal{I}\equiv\sqrt{\frac{3\epsilon_{A}}{2\epsilon_{\phi}}},$ (4)
-which characterizes the ratio between two rolling velocity of the fields.
-[…]
+If the energy density of the vector field remains subdominated compared with inflaton $\rho_{A}\ll V$, the anisotropy of spacetime is always of the order of slow-roll parameters [Watanabe:2010fh], i.e., $aH\simeq(-\eta)^{-1-\epsilon_{H}}$ and $bH\simeq(-\eta)^{-1-\epsilon_{H}-\epsilon_{A}}$ so that $b^{\prime}/b-a^{\prime}/a\simeq aH\epsilon_{A}$, where $H\equiv\dot{a}/a$ and the prime denotes the derivative with respect to conformal time $\eta$.
+The slow-roll parameters are defined by $\epsilon_{\phi}\equiv\dot{\phi}^{2}/2M_{\text{pl}}^{2}H^{2},\epsilon_{A}\equiv f^{2}\dot{A}^{2}/3a^{2}M_{\text{pl}}^{2}H^{2}$ in our ansatz.
+During the anisotropic phase, the potential of inflaton still dominates the inflation, while a portion of its energy transforms to the vector field.
+In this scenario, both the rolling of the inflaton and the vector field are made responsible for the deviation of the de Sitter space, i.e., $\epsilon_{H}=\epsilon_{\phi}+\epsilon_{A}$.
+A concrete realization of ansiotropic inflation can be achieved by choosing the kinetic coupling as $f(\phi)=\exp\big(2c\int d\phi\ V/V_{\phi}M_{\text{pl}}^{2}\big)$ with $c>1$, where $V_{\phi}\equiv dV/d\phi$ [Watanabe:2009ct, Maleknejad:2012fw].
+Inflation can transition into an anisotropic attractor with a constant background electric field [Watanabe:2009ct, Maleknejad:2012fw, Kanno:2010nr, Chen:2022ccf].
+The inflaton rolling velocity is then reduced to $\epsilon_{\phi}\simeq\epsilon_{\phi}/c$, and the energy density of the vector field is given by $\rho_{E}/V\equiv f^{2}\dot{A}^{2}/2a^{2}\simeq\epsilon_{H}(c-1)/2c$.
+The parameter $\mathcal{I}$ is fundamentally unconstrained in this background dynamics, with its range $0<\mathcal{I}<\infty$ determined by $\mathcal{I}\simeq\sqrt{3(c-1)/2}$ in the anisotropic phase.
+Our analysis will focus particularly on the regime where $\mathcal{I}\gg 1$.
 III Strong mixing of perturbations
-The vector field breaks the 3-d rotational symmetry of the universe to a 2-d rotational one in the $yz$ plane.
-This symmetry breaking complicates the metric perturbations, with the complete quadratic action having been thoroughly analyzed in previous works [22, 23, 24, 25, 26, 27, 28].
-However, it is known that the dominated contributions to the curvature perturbation arise only from matter perturbations in the spatially-flat gauge, while the interactions between matter and metric are slow-roll suppressed.
+The vector field breaks the 3D rotational symmetry of the Universe to a 2D rotational one in the $yz$ plane.
+This symmetry breaking complicates the metric perturbations, with the complete quadratic action having been thoroughly analyzed in previous works [Watanabe:2010fh, Emami:2013bk, Bartolo:2012sd, Gumrukcuoglu:2007bx, Himmetoglu:2009mk, Gumrukcuoglu:2010yc, Durrer:2024ibi].
+However, it is known that the dominated contributions to the curvature perturbation arise only from matter perturbations in the spatially flat gauge, while the interactions between matter and metric are slow-roll suppressed.
 Consequently, we may safely neglect gravitational back-reaction in the following.
@@
 $\phi=\phi_{0}+\varphi,\ \ \ \ \ A_{\mu}^{(S)}=(\delta A_{0},A+\delta A_{1},0,0).$ (5)
-Here we consider only the scalar parts of the perturbations of $A_{\mu}$ and have removed d.o.f by fixing the U(1) symmetry.
+Here, we consider only the scalar parts of the perturbations of $A_{\mu}$ and have removed d.o.f.
+by fixing the U(1) symmetry.
 Without loss of generality, we align the wave number as $\bm{k}=(k_{1},k_{2},0)$, where $k_{1}=k\cos\theta$ and $k_{2}=(b/a)k\sin\theta$, and $\theta$ being the angle between the vector field and $k$.
-After eliminating the non-dynamical d.o.f.
+After eliminating the nondynamical d.o.f.
 $\delta A_{0}$ and introducing a new variable $\mathcal{A}\equiv(f/a)\sin\theta\delta A_{1}$, the dominant quadratic action is given by
@@
 $-12H^{2}\mathcal{I}\sin\theta\Big(\varphi^{*}\mathcal{A}+c.c.\Big)\Bigg].$ (6)
-Here, we have assumed $\dot{\phi}<0$, $\dot{A}>0$, and treat $\mathcal{I}$ as a constant to leading-order of the slow-roll approximation.
-Spacetime anisotropy, being slow-roll suppressed, is also neglected.
+Here, we have assumed $\dot{\phi}<0$, $\dot{A}>0$, and treat $\mathcal{I}$ as a constant to leading order of the slow-roll approximation.
+Spacetime anisotropy, being slow-roll suppressed, is also neglected so that $a=b$.
 The last two terms in (III) represent the mixing between $\varphi$ and $\mathcal{A}$, which is controlled by $\mathcal{I}\sin\theta$.
 We are interested in the strong mixing $\mathcal{I}\sin\theta\gg 1$, which is the subject of the rest of this paper.
-One can immediately note that in this regime the mass $m^{2}_{\varphi}\equiv 8\mathcal{I}^{2}H^{2}\cos{2\theta}$ can become heavy for some angle values $\theta$.
-If this is the case, there is a hierarchy between two frequencies $\omega_{\pm}^{2}=p^{2}+4\mathcal{I}^{2}H^{2}\pm 4\mathcal{I}H(\mathcal{I}^{2}H^{2}+p^{2}\sin^{2}\theta)^{1/2}$ of modes on sub-horizon scales [40].
-When the momentum of the modes redshifts to $p\ll|m_{\varphi}|$, the high-frequency mode $\omega_{+}$ can be integrated out, yielding a low-energy effective theory for the light modes [41, 42, 43].
-Specifically, To obtain the leading-order solution for $\varphi$, we expand the propagator $1/(m^{2}_{\varphi}-\square)=1/m^{2}_{\varphi}+\square/m^{4}_{\varphi}+\cdots$ in the equation of motion.
+One can immediately note that in this regime, the mass $m^{2}_{\varphi}\equiv 8\mathcal{I}^{2}H^{2}\cos{2\theta}$ can become heavy for some angle values $\theta$.
+If this is the case, there is a hierarchy between two frequencies $\omega_{\pm}^{2}=p^{2}+4\mathcal{I}^{2}H^{2}\pm 4\mathcal{I}H(\mathcal{I}^{2}H^{2}+p^{2}\sin^{2}\theta)^{1/2}$ of modes on subhorizon scales [Achucarro:2012yr].
+Here, $p\equiv k/a$ is the physical momentum.
+When the momentum of the modes redshifts to $p\ll|m_{\varphi}|$, the high-frequency mode $\omega_{+}$ can be integrated out, yielding a low-energy effective theory for the light modes [Cheung:2007st, Baumann:2011su, Gwyn:2012mw].
+Specifically, to obtain the leading-order solution for $\varphi$, we denote operator $\square\equiv-\partial^{2}/\partial t^{2}-3H\partial/\partial t-k^{2}/a^{2}$ and expand the propagator $1/(m^{2}_{\varphi}-\square)=1/m^{2}_{\varphi}+\square/m^{4}_{\varphi}+\cdots$ in the equation of motion.
 Then we can solve this leading solution as
@@
 This is an effective action of a massless perturbation with sound speed $c_{s}$.
-Remarkably, $c_{s}^{2}$ becomes imaginary for $\theta\in(\pi/4,3\pi/4)\cup(5\pi/4,7\pi/4)$, leading to exponentially growing solutions before the horizon crossing  [^3] Recently, the multi-field inflation with strongly non-geodesic attractors in the field space has been extensively explored [57, 58].
-The effective theory of inflationary dynamics is a massless mode with imaginary sound speed [30, 31, 32, 33, 34, 35].
-It can also provide the mechanism for generating PBH in the multi-filed scenario [36, 37].
-Consequently, the power spectrum of $\mathcal{A}$ is scale-invariant(of leading order) and enhanced on large scales.
+Remarkably, $c_{s}$ becomes imaginary for $\theta\in(\pi/4,3\pi/4)\cup(5\pi/4,7\pi/4)$, leading to exponentially growing solutions before the horizon crossing  [^3] Recently, the multi-field inflation with strongly non-geodesic attractors in the field space has been extensively explored [Brown:2017osf, Bjorkmo:2019fls].
+The effective theory of inflationary dynamics is a massless mode with imaginary sound speed [Garcia-Saenz:2018ifx, Fumagalli:2019noh, Bjorkmo:2019qno, Garcia-Saenz:2019njm, Garcia-Saenz:2018vqf, Garcia-Saenz:2025jis].
+It can also provide the mechanism for generating PBH in the multi-filed scenario [Palma:2020ejf, Fumagalli:2020adf].
+Consequently, the power spectrum of $\mathcal{A}$ is scale invariant(of leading order) and enhanced on large scales.
 This enhancement is anisotropic due to the angle-dependent sound speed $c_{s}(\theta)$.
 On the other hand, if $\sin\theta\simeq 0$, $\varphi$ and $\mathcal{A}$ are decoupled and $\varphi$ is also heavy with a real mass and decay away on the superhorizon for any value of large $\mathcal{I}$.
-The enhanced solution for $\mathcal{A}$ on the superhorizon can be found by solving the coupled system inside the horizon [32, 45, 46].
-Here we are only interested in the enhanced solution, where $\sin^{2}{\theta}>1/2$.
+The enhanced solution for $\mathcal{A}$ on the superhorizon can be found by solving the coupled system inside the horizon [Bjorkmo:2019qno, Christodoulidis:2023eiw, Chen:2023bcz].
+Here, we are only interested in the enhanced solution, where $\sin^{2}{\theta}>1/2$.
 The conjugate momentum of the perturbations is given by
@@
 $J_{1}=$ $\left(\ \begin{matrix}0&0&0&0&\\ 0&-3&-12\mathcal{I}\sin{\theta}&0&\\ 0&0&0&0&\\ -12\mathcal{I}\sin{\theta}&0&0&-3&\end{matrix}\right).$ (11)
-Considering the Hubble term as a small correction to the system on subhorizon, the matrix $J_{0}$ is of 00-$th$ order while $J_{1}$ is $1$-$st$ order.
-Near the horizon crossing, the exponential growth mode of $\mathcal{A}$ emerges from the imaginary frequency $\omega_{-}$ of the coupled system when $p<|m_{\varphi}|$, which yields the enhanced solution at the limit $-k\eta\to 0$ [47],
+Considering the Hubble term as a small correction to the system on subhorizon, the matrix $J_{0}$ is of $0th$ order while $J_{1}$ is first order.
+Near the horizon crossing, the exponential growth mode of $\mathcal{A}$ emerges from the imaginary frequency $\omega_{-}$ of the coupled system when $p<|m_{\varphi}|$, which yields the enhanced solution at the limit $k\eta\to 0$ [supp],
 $\mathcal{A}\simeq\left|2-4\sin^{2}\theta\right|^{1/4}e^{2\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}.$ (12)
@@
 The PBH can be formed if the curvature perturbation generated on small scales during inflation is enhanced.
-For an anisotropic inflationary universe, the uniform-energy curvature perturbation we can use in the spatially-flat gauge is
+For an anisotropic inflationary universe, the uniform-energy curvature perturbation we can use in the spatially flat gauge is
 $\zeta\simeq-H\frac{\delta\rho}{\dot{\rho}}.$ (14)
@@
 $\mathcal{R}=H\frac{ik^{a}}{k^{2}}V_{a}$ (15)
-in the spatially-flat gauge, where $a=1,2,3$ and the total velocity $V_{a}\equiv(\rho+P)^{-1}T^{t}_{\ a}$.
+in the spatially flat gauge, where $a=1,2,3$ and the total velocity $V_{a}\equiv(\rho+P)^{-1}T^{t}_{\ a}$.
 This reduces to the standard comoving curvature perturbation $\mathcal{R}_{\text{iso}}=-H(\rho+P)u$ in isotropic backgrounds, where $u$ is the scalar velocity of fluids $v_{a}=\partial_{a}u+u_{a}$.
-One can find that in our configuration of anisotropic inflation, $\mathcal{R}=-\zeta$ holds on superhorizon scales up to slow-roll corrections from spacetime anisotropy [47].
+One can find that in our configuration of anisotropic inflation, $\mathcal{R}=-\zeta$ holds on superhorizon scales up to slow-roll corrections from spacetime anisotropy [supp].
 We therefore use $\mathcal{R}$ to compute the power spectrum, consistent with our treatment of neglecting slow-roll suppressed terms.
@@
 $\mathcal{R}=\frac{1}{\sqrt{2\epsilon_{H}}M_{\text{pl}}H}\frac{-\varphi+\mathcal{I}\sin\theta\mathcal{A}}{\sqrt{1+2\mathcal{I}^{2}/3}}.$ (16)
-In the EFT theory of $\mathcal{A}$ (13), the dominant contribution to the curvature perturbation is from the perturbation of the vector field $\mathcal{A}$ due to $\mathcal{R}_{\mathcal{A}}/\mathcal{R}_{\varphi}\sim\mathcal{I}^{2}\gg 1$.
+In the EFT theory of $\mathcal{A}$ (13), the dominant contribution to the curvature perturbation is from the perturbation of the vector field $\mathcal{A}$ due to the contributions in (16) of two perturbations $\sim\mathcal{I}^{2}\gg 1$.
 If we assume that $\epsilon_{H}$ is a constant during inflation, we have
 $\mathcal{R}\propto\sqrt{\frac{3}{2}}\sin{\theta}\mathcal{A}.$ (17)
-The Fourier modes of $\mathcal{R}$ can be decomposed in terms of two sets of operators since we have two perturbations: ~$\tilde{\mathcal{R}}(\bm{k},\eta)=\sum_{\alpha=\pm}\left[\mathcal{R_{\alpha}}(k,\eta)a_{\alpha}(\bm{k})+\mathcal{R^{*}_{\alpha}}(k,\eta)a_{\alpha}^{\dagger}(-\bm{k})\right]$, where $a_{\alpha}^{\dagger}(\bm{k})$ and $a_{\alpha}(\bm{k})$ are creation and annihilation operators that satisfy the commutation relation $[a_{\alpha}(\bm{k}),a_{\beta}^{\dagger}(\bm{k^{\prime}})]=(2\pi)^{3}\delta_{\alpha\beta}\delta^{(3)}(\bm{k}-\bm{k^{\prime}})$.
+The Fourier modes of $\mathcal{R}$ can be decomposed in terms of two sets of operators since we have two perturbations: $\tilde{\mathcal{R}}(\bm{k},\eta)=\sum_{\alpha=\pm}\left[\mathcal{R_{\alpha}}(k,\eta)a_{\alpha}(\bm{k})+\mathcal{R^{*}_{\alpha}}(k,\eta)a_{\alpha}^{\dagger}(-\bm{k})\right]$, where $a_{\alpha}^{\dagger}(\bm{k})$ and $a_{\alpha}(\bm{k})$ are creation and annihilation operators that satisfy the commutation relation $a_{\alpha}(\bm{k}),a_{\beta}^{\dagger}(\bm{k^{\prime}})]=(2\pi)^{3}\delta_{\alpha\beta}\delta^{(3)}(\bm{k}-\bm{k^{\prime}})$.
 The dimensionless power spectrum of the curvature perturbations is then computed by
 $\mathcal{P}_{\mathcal{R}}(k)=\frac{k^{3}}{2\pi^{2}}\sum_{\alpha=\pm}|\mathcal{R}_{\alpha}(k)|^{2}.$ (18)
-Here we consider a transient anisotropic phase, i.e., $\mathcal{I}=\mathcal{I}_{m}\gg 1$ in a period $\Delta N\gtrsim\log\mathcal{I}$ near $N_{f}>N_{\text{CMB}}$, to see how a time-dependent $\mathcal{I}(\eta)$ can result in an anisotropic peak of the curvature power spectrum on small scales  [^4] We assume that at time $N\ll N_{f}$ and $N\gg N_{f}$, inflation is in the anisotropic phase with $\mathcal{I}\ll 1$ so that the gauge field $\mathcal{A}$ is approximately a massless field..
-In this case, only modes with wavenumber $k\sim k_{f}\equiv a(N_{f})H$ can experience exponential growth when crossing the horizon.
+Here, we consider a transient anisotropic phase, i.e., $\mathcal{I}\gg 1$, in a period $\Delta N\gtrsim\log\mathcal{I}$.
+We assume that $\mathcal{I}$ reaches its maximum value $\mathcal{I}(N_{f})=\mathcal{I}_{\text{max}}$ at time $N_{f}>N_{\text{CMB}}$, to see how a time-dependent $\mathcal{I}(\eta)$ can result in an anisotropic peak of the curvature power spectrum on small scales  [^4] We assume that at time $N\ll N_{f}$ and $N\gg N_{f}$, inflation is in the anisotropic phase with $\mathcal{I}\ll 1$ so that the vector field $\mathcal{A}$ is approximately a massless field..
+The momentum corresponding to modes that cross the horizon at time $N_{f}$ is defined as $k_{f}\equiv a(N_{f})H$.
+In this case, only modes with wave number $k\sim k_{f}$ can experience exponential growth when crossing the horizon.
 The power spectrum of the enhanced curvature perturbation near $k\sim k_{f}$ can be estimated as
@@
 Even a slight deviation $\Delta\theta\sim\mathcal{I}^{-1}$ from $\pi/2$ leads to a drastic suppression of the peak of the power spectrum.
-On the other hand, we define the mass-crossing time $k/a(\tilde{N})=|m_{\varphi}(\tilde{N})|$ for any $k$ mode.
-Because most of the growth occurs at the early time [32, 45, 37].
-The location of the peak $k_{p}$ is determined by $\tilde{N}\sim N_{f}$, i.e.
-the mode that crosses the mass-horizon at time $N_{f}$,
+[…]
 $k_{p}(\theta)\sim\frac{|m_{\varphi}(\theta)|}{H}\cdot k_{f},$ (20)
 which means that the location of the peaks is also anisotropic.
-Modes with $k\gg k_{f}$ experience nothing because the vector field has been turned off well before horizon-crossing.
-On the other hand, large-scale modes $k\gg k_{f}$ are still almost scale-invariant but are kicked by the vector, which results in a shift deformation.
+Modes with $k\gg k_{f}$ experience nothing because the vector field has been turned off well before horizon crossing.
+On the other hand, large-scale modes $k\gg k_{f}$ are still almost scale invariant but are kicked by the vector, which results in a shift deformation.
 The amplitudes of the shift obviously depend on the angle $\theta$, and the large-scale power spectrum is significantly anisotropy.
@@
 In the early stage of inflation, the energy density of $A$ is negligible, reducing the system to a two-field system.
-Assuming slow-roll conditions for both $\psi$ and $\phi$, adiabatic and entropy perturbations remain decoupled at early times [49].
+Assuming slow-roll conditions for both $\psi$ and $\phi$, adiabatic and entropy perturbations remain decoupled at early times [Gordon:2000hv].
 Consequently, the curvature perturbation remains conserved on large scales.
@@
 Figure 1: The power spectrum of $\mathcal{R}$ for our toy example (IV.2).
-We use the Gaussian profile $\mathcal{I}(N)=\mathcal{I}_{\text{max}}e^{-(N-N_{f})^{2}/(2\Delta N^{2})}$ with $\mathcal{I}_{\text{max}}=9$, $\Delta N=2$, $N_{f}\simeq N_{\text{CMB}}+5$ and $r_{\text{ini}}=10$.
+We use the Gaussian profile $\mathcal{I}(N)=\mathcal{I}_{\text{max}}e^{-(N-N_{f})^{2}/(2\Delta N^{2})}$ with $\mathcal{I}_{\text{max}}=5$, $\Delta N=2$, $N_{f}\simeq N_{\text{CMB}}+5$ and $r_{\text{ini}}=10$.
+$\mathcal{P}_{\mathcal{A}}=\mathcal{P}_{0}\left|2-4\sin^{2}\theta\right|^{1/2}e^{4\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}\big|_{N_{v}(k)}$, where $N_{v}(k)$ is given by $p(N_{v})=|m_{\varphi}(N_{v})|$.
 The black curve represent peaks $k_{p}(\theta)$ of $\mathcal{P}_{\mathcal{R}}$.
@@
 One may be concerned that the exponential enhancement of perturbations leads to an incredible enhancement on NG.
-However, it has been shown that the low-energy EFT with imaginary sound speed reduced from strongly non-geodesic attractors of multifield inflation generates only mild NG [32, 31, 33, 34].
-Moreover, the perturbative validity is also well controlled in this attractor [31, 32, 35].
-We believe that this analysis can be used in our case in this Letter.
+However, it has been shown that the low-energy EFT with imaginary sound speed reduced from strongly nongeodesic attractors of multifield inflation generates only mild NG [Bjorkmo:2019qno, Fumagalli:2019noh, Garcia-Saenz:2019njm, Garcia-Saenz:2018vqf].
+Moreover, the perturbative validity is also well controlled in this attractor [Fumagalli:2019noh, Bjorkmo:2019qno, Garcia-Saenz:2025jis].
+We believe that this analysis can be used in our case in this paper.
+Moreover, we only consider $\Delta N\gtrsim\log\mathcal{I}$ in this paper.
+References [Palma:2020ejf, Fumagalli:2020adf] demonstrate that the curvature perturbation power spectrum exhibits oscillatory features when $\Delta N\lesssim\log\mathcal{I}$.
+It is interesting to explore this case under our anisotropic model.
+The anisotropic hair may imprint distinctive features in this power spectrum.
 The anisotropic features of the peak can be caught by some future observations, such as PBH production.
@@
 It is necessary to figure out the physical details of this case.
-In particular, anisotropic hair may be imprinted on the initial spin of PBHs, leaving behind intriguing observational signatures [50].
+In particular, anisotropic hair may be imprinted on the initial spin of PBHs, leaving behind intriguing observational signatures [Harada:2024jxl].
 Moreover, enhancements of the curvature power spectrum on small scales are generally accompanied by the production of scalar-induced gravitational waves(SIGWs).
 This implies that anisotropic features will also manifest as sources in the SIGW spectrum.
-The frequencies of SIGWs is related to $N_{f}$ as $N_{f}+\log\left(\mathcal{I}\right)\sim\log\left(f_{\text{peak}}/\text{Hz}\right)+37$ [51], which spans a broad spectrum covered by gravitational-wave experiments, such as PTA and LISA.
-A model-independent discussion of SIGWs sourced from a statistically anisotropic scalar spectrum has been studied in [52, 53].
-We provide a specific realization of a statistically anisotropic peak in this Letter.
-It is interesting to explore signatures of this anisotropic hair in the SIGWs and test it in future observations.
+The frequencies of SIGWs is related to $N_{f}$ as $N_{f}+\log\left(\mathcal{I}\right)\sim\log\left(f_{\text{peak}}/\text{Hz}\right)+37$ [Fumagalli:2020nvq], which spans a broad spectrum covered by gravitational-wave experiments, such as PTA and LISA.
+A model-independent discussion of SIGWs sourced from a statistically anisotropic scalar spectrum has been studied in [Chen:2022qec, Kuang:2023urj].
+We provide a specific realization of a statistically anisotropic peak in this paper.
+Moreover, the scalar modes can be the source of GWs generated during inflation.
+The GWs are enhanced by the presence of a scalar excited state and overcome the SIGWs [Fumagalli:2021mpc].
+It is interesting to explore signatures of this anisotropic hair in the stochastic gravitational wave background(SGWB) and test it in future observations.
 Acknowledgements.